OpenStudy (unklerhaukus):

\[\newcommand\dd[1]{\,\mathrm d#1}% infinitesimal \newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd #4}}% integral _{a}^{b}{f(x)}\dd x \newcommand\erf[1]{\operatorname {erf}\left(#1\right)}% Error function (x) \newcommand\erfi[1]{\frac2{\sqrt \pi}\intl{0}{#1}{e^{-u^2}}{u}}% Error function integral (x) {-------------------}\\ \text{From the definition prove the result:} \\ \erf x=\frac2{\sqrt\pi}\left\{x-\frac{x^3}3+\frac{x^5}{2!5}\color{brown}-\frac{x^7}{3!7}+\dots\right\}\\ {-------------------}\]

OpenStudy (unklerhaukus):

\[\begin{align*} \\ \erf x &=\erfi x\\ &=\\ &\,\vdots\\ &=\\ &=\frac2{\sqrt\pi}\left\{x-\frac{x^3}3+\frac{x^5}{2!5}+\frac{x^7}{3!7}+\dots\right\} \end{align*}\]

OpenStudy (unklerhaukus):

I'm not sure how to turn the integral into the infinite series,

OpenStudy (aravindg):

as always this is above my level always wish i could help !

Parth (parthkohli):

Looks like a Maclaurin Series bro.

OpenStudy (unklerhaukus):

are you series ?

Parth (parthkohli):

lol\[f(x) = \sum_{n = 0}^{\infty}\dfrac{{f^n}(0)}{n!}x^n\]Definitely not a Maclaurin.

Parth (parthkohli):

Not sure. I am not good at this :-(

OpenStudy (unklerhaukus):

ah yes Maclaurin Series (13) http://mathworld.wolfram.com/MaclaurinSeries.html

OpenStudy (anonymous):

looks like the series is \[ \frac{x^{2n+1}}{n!(2n+1)} \]

OpenStudy (anonymous):

But that - sign in there is throwing me off

OpenStudy (unklerhaukus):

\[\newcommand\dd[1]{\,\mathrm d#1}% infinitesimal \newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd #4}}% integral _{a}^{b}{f(x)}\dd x \newcommand\erf[1]{\operatorname {erf}\left(#1\right)}% Error function (x) \newcommand\erfi[1]{\frac2{\sqrt \pi}\intl{0}{#1}{e^{-u^2}}{u}}% Error function integral (x) \newcommand\de[2]{\frac{\mathrm d #1}{\mathrm d#2}} % first order derivative \erf x =\erfi x \\ \de{\erf x}x =\de{}x\erfi x\\ \qquad\quad\quad=\frac2{\sqrt\pi}e^{-x^2} \]

OpenStudy (anonymous):

Why not use: \[ \Large e^{x} = \sum_{n=0}^{\infty } \frac{x^n}{n!} \implies e^{-x^2} = \sum_{n=0}^{\infty } \frac{(-x^2)^n}{n!} \]

OpenStudy (anonymous):

It's really easy to find the anti derivative

OpenStudy (unklerhaukus):

i think i got it now ! thanks!

OpenStudy (anonymous):

No problem,

OpenStudy (anonymous):

the part that threw me off was that there wasn't an alternating - +

OpenStudy (unklerhaukus):

oh dear i made an error in the question , it should be alternating , like you say @wio

OpenStudy (anonymous):

@UnkleRhaukus Where are you getting the questions from anyway?

Parth (parthkohli):

What is all this `\newcommand` thing? Are we allowed to introduce new variables in MathJax?

OpenStudy (anonymous):

Is it a class you're getting it from?

OpenStudy (anonymous):

Or a book or what?

OpenStudy (anonymous):

@ParthKohli you're allowed to make command shortcuts, but I think it only works for your own post.

OpenStudy (unklerhaukus):

MATH202 Differential equations sophomore subject \[\emph{Exercise 4D}1(d)\]

OpenStudy (unklerhaukus):

Are new commands allowed Ambassador?

OpenStudy (experimentx):

don't expand that series ... just integrate inside summation sign.

OpenStudy (experimentx):

there are many cases where you have to change functions into infinite series to evaluate integral with real methods. I think ... expanding the series will just mess up.

OpenStudy (unklerhaukus):

have i done something wrong?

OpenStudy (experimentx):

no it's okay ... just advised for simplicity.